OpenStudy (anonymous):

A car's gas tank held 22 gallons of gas. Six days later, 7 gallons of gas remained in the tank. Find the average rate of change of gas per day. Show work.

OpenStudy (anonymous):

(22-7)/6 = 15/6 = 5/2 gallons of gas/day

OpenStudy (anonymous):

thank you